# Computing Lower and Upper Hitting Probabilities for Imprecise Markov Chains

### v1

@@ 3.3 Upper Hitting Probabilities
Let us start with the claims about the structure of $\overline{p}^{\mathcal{T}}$. We already know that Equation (17) has a minimal nonnegative solution $\overline{p}^{\mathcal{T}}\in[0,1]^{\mathcal{X}}$ [krak2019hitting, Corollary 19]. By the definition of the system, this solution satisfies $\overline{p}^{\mathcal{T}}(x)=1$ for $x\in A$. The fact that $\overline{p}^{\mathcal{T}}(x)=0$ if and only if $x\in{\mathscr{W}_{\mathcal{T}}}$ follows from the definition of upper hitting probability and Lemma 3.11.
Now we prove the uniqueness of $\overline{p}^{\mathcal{T}}$. By the minimality of $\overline{p}^{\mathcal{T}}$ we have $\overline{p}^{\mathcal{T}}\leq p$ pointwise, for all $p\in[0,1]^{\mathcal{X}}$ solution to the system in Equation (17) with $p(x)=1$ for all $x\in A$ and $p(x)=0$ for all $x\in{\mathscr{W}_{\mathcal{T}}}$. Set
$d\coloneqq p-\overline{p}^{\mathcal{T}}\geq 0.$
Subtracting the two fixed-point equations yields
$d=\mathbb{1}_{A^{c}}\cdot\big(\overline{T}p-\overline{T}\overline{p}^{\mathcal{T}}\big).$
Since $p=\overline{p}^{\mathcal{T}}+d$ and $d\geq 0$, subadditivity (Definition 2.15(2)) gives
$\overline{T}p=\overline{T}(\overline{p}^{\mathcal{T}}+d)\leq\overline{T}\overline{p}^{\mathcal{T}}+\overline{T}d.$
Hence $\overline{T}p-\overline{T}\overline{p}^{\mathcal{T}}\leq\overline{T}d,$ and therefore
$d=\mathbb{1}_{A^{c}}\cdot\big(\overline{T}p-\overline{T}\overline{p}^{\mathcal{T}}\big)\leq\mathbb{1}_{A^{c}}\cdot\overline{T}d.$
Applying this inequality $n$ times yields
$d\leq\mathbb{1}_{A^{c}}\cdot\overline{T}\left(\mathbb{1}_{A^{c}}\cdot\overline{T}d\right)\leq\mathbb{1}_{A^{c}}\cdot\overline{T}\left(\,\mathbb{1}_{A^{c}}\cdot\overline{T}\left(\dots\left(\mathbb{1}_{A^{c}}\cdot\overline{T}d\right)\right)\right),$
which we write
$d\leq\left(\mathbb{1}_{A^{c}}\cdot\overline{T}\right)^{n}d.$
[…]
for all $n\in\mathbb{N}$. Therefore, we have
[…]
for all $n\in\mathbb{N}$.
Moreover, we observe that
$\Big(\left(\mathbb{1}_{A^{c}}\cdot\overline{T}\right)^{n}\mathbb{1}_{A^{c}}\Big)\Big|_{A^{c}}=\left(\overline{T}|_{A^{c}}\right)^{n}\mathbb{1}_{A^{c}}.$
For each $x\in A^{c}\setminus\mathscr{W}_{T}$ the quantity $((\overline{T}|_{A^{c}})^{n}\mathbb{1}_{A^{c}})(x)$ is the upper probability of avoiding $A$ for $n$ steps and therefore decays to 0 as $n$ approaches infinity. Taking the limit on both sides of (18) yields $d(x)\leq 0$ for every $x\in A^{c}\setminus\mathscr{W}_{\mathcal{T}}$, so $d=0$, which is what we wanted to prove. ∎
@@ 3.4 Further Discussions on the Definition of Reachability
Proposition 3.9 shows that this definition is equivalent to $\operatorname{(LR0)}$: indeed, $x\in\mathcal{A}_{\mathcal{T}}$, i.e. $x\not\rightharpoondown A$, if and only if $\underline{p}^{\mathcal{T}}(x)=0$. Taking the negation, we obtain the equivalence between $\operatorname{(LR0)}$ and $\operatorname{(LRHP)}$.
In the same way, the definition of upper reachability can be rewritten in terms of upper hitting probabilities.
@@ 4.1 Computing Lower Hitting Probabilities
We show an algorithmic result for computing $\underline{p}^{\mathcal{T}}$. This result, which emulates Krak [krak2021comphit, Proposition 7], provides a practical algorithm for calculating $\underline{p}^{\mathcal{T}}$ iterating over extreme points of $\mathcal{T}$. The only condition is to start with a $T_{1}\in\mathcal{T}$ such that the trivial hitting probabilities are already correct (i.e. $p_{1}(x)=0$ for $x\in\mathcal{A}_{\mathcal{T}}$ and $p_{1}(x)=1$ for $x\in A$). Definition (LR1) provides a practical algorithm to identify the set $\mathcal{A}_{\mathcal{T}}$. In particular, if we start with $\mathcal{D}=A$ then $\mathcal{D}_{k^{*}}^{c}=\mathcal{A}_{\mathcal{T}}$. The matrix $T\in\mathcal{T}$ found in the last iteration of this algorithm is the desired starting matrix.
@@
Let $T_{1}\in\mathcal{T}$ be any transition matrix such that $\mathcal{C}_{T_{1}}=\mathcal{A}_{\mathcal{T}}$. For all $n\in\mathbb{N}$, let $p_{n}$ be the minimal nonnegative solution in $0,1]^{\mathcal{X}}$ to the linear system
$p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n}p_{n},$
@@
Then, the sequence $(p_{n})_{n\in\mathbb{N}}$ is nonincreasing and its limit $p^{*}\coloneqq\lim_{n\to+\infty}p_{n}$ is the unique minimal solution of the system
$\underline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot\underline{T}\,\underline{p}^{\mathcal{T}}.$ (19)
@@
The proof uses induction on (20). For the base case $n=1$, we have $\mathcal{C}_{T_{1}}=\mathcal{A}_{\mathcal{T}}$ by the choice of $T_{1}\in\mathcal{T}$. Assume now that $\mathcal{C}_{T_{n}}=\mathcal{A}_{\mathcal{T}}$ for some $n\in\mathbb{N}$. We will show that also $\mathcal{C}_{T_{n+1}}=\mathcal{A}_{\mathcal{T}}$. Let $x\in\mathcal{A}_{\mathcal{T}}$, then
$T_{n}p_{n}](x)=\sum_{y\in\mathcal{X}}T_{n}(x,y)p_{n}(y)=p_{n}(x)=0,$
since $\mathcal{A}_{\mathcal{T}}\subseteq A^{c}$ and $p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n}p_{n}$. Note that the last equality holds by our induction hypothesis in combination with Proposition 2.5. By construction, we have
$T_{n+1}p_{n}=\underline{T}p_{n}\leq T_{n}p_{n},$
since $T_{n}\in\mathcal{T}$. Therefore,
$0=[T_{n}p_{n}](x)\geq[T_{n+1}p_{n}](x)=\sum_{y\in\mathcal{X}}T_{n+1}(x,y)p_{n}(y).$
Since the right-hand side is nonnegative, we must have $T_{n+1}(x,y)=0$ for every $y\in\mathcal{X}$ with $p_{n}(y)>0$. By Proposition 2.5 this implies $T_{n+1}(x,y)=0$ for all $y\notin\mathcal{C}_{T_{n}}$. Applying the induction hypothesis, we obtain $T_{n+1}(x,y)=0$ for every $y\in\mathcal{A}_{\mathcal{T}}^{c}$, and hence $x\not\to^{T_{n+1}}y$ for all $y\in\mathcal{A}_{\mathcal{T}}^{c}$. Since $\mathcal{A}_{\mathcal{T}}\subseteq A^{c}$, it follows that $x\in\mathcal{C}_{T_{n+1}}$. As $x\in\mathcal{A}_{\mathcal{T}}$ was arbitrary, $\mathcal{A}_{\mathcal{T}}\subseteq\mathcal{C}_{T_{n+1}}$. Finally, using Lemma 3.8, which states $\mathcal{A}_{\mathcal{T}}=\bigcup_{T\in\mathcal{T}}\mathcal{C}_{T}$, we conclude that $\mathcal{A}_{\mathcal{T}}=\mathcal{C}_{T_{n+1}}$, completing the induction. ∎
@@
Let us first establish that the sequence $(p_{n})_{n\in\mathbb{N}}$ is well-defined. Since each $p_{n}$ is a vector of hitting probabilities, we have that $p_{n}\in[0,1]^{{\mathcal{X}}}$ for all $n\in\mathbb{N}$. Moreover, for any $n\in\mathbb{N}$, the minimal nonnegative solution in $0,1]^{{\mathcal{X}}}$ to
$p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n}p_{n}$
exists due to Proposition 2.5. Furthermore, for any $n\in\mathbb{N}$, the existence of the extreme point $T_{n+1}$ is guaranteed by Proposition 2.18.
Next, we will show that $(p_{n})_{n\in\mathbb{N}}$ is convergent and nonincreasing. By Proposition 2.5, we have that $p_{n}(x)=1$ for $x\in A$ for all $n\in\mathbb{N}$. Therefore $(p_{n})_{n\in\mathbb{N}}$ is trivially nonincreasing and convergent on $A$. By Lemma 4.2, we have that $\mathcal{A}_{\mathcal{T}}=\mathcal{C}_{T_{n}}$ for all $n\in\mathbb{N}$, therefore by Proposition 2.5 it follows that $p_{n}(x)=0$ for all $x\in\mathcal{A}_{\mathcal{T}}$ and all $n\in\mathbb{N}$, meaning that $(p_{n})_{n\in\mathbb{N}}$ is trivially nonincreasing and convergent on $\mathcal{A}_{\mathcal{T}}$.
Next, we will show that $(p_{n})_{n\in\mathbb{N}}$ is nonincreasing on $A^{c}\setminus\mathcal{A}_{\mathcal{T}}$ as well. This is shown analogously to the proof of Krak’s Proposition 7 [krak2021comphit]. Consider any $n\in\mathbb{N}$. Recall that
$T_{n+1}p_{n}=\underline{T}p_{n}\leq T_{n}p_{n}.$
This then yields that
@@
where we used Lemma 2.6 for the last step in combination with the fact that we have already shown that $\mathcal{A}_{\mathcal{T}}=\mathcal{C}_{T_{n}}$ for all $n\in\mathbb{N}$. Repeatedly expanding this inequality $m\in\mathbb{N}$ times yields that
@@
Letting $m\to+\infty$ and applying Lemma 2.12 then yields that
@@
where the last step follows from Proposition 2.13. Thus the sequence $(p_{n})_{n\in\mathbb{N}}$ is nonincreasing. Since $(p_{n})_{n\in\mathbb{N}}$ is nonincreasing and $p_{n}\in[0,1]^{{\mathcal{X}}}$ for all $n\in\mathbb{N}$, it follows that the limit $p^{*}\coloneqq\lim_{n\to+\infty}p_{n}$ exists and lives in $0,1]^{{\mathcal{X}}}$.
Let us move on to showing that $p^{*}$ satisfies the system in Equation (19). Fix any $n\in\mathbb{N}$. Then
@@
Continuing (21), we see that
@@
where we used Lemma 2.16(3) for the second and third inequalities. Since all summands on the final right-hand side of (22) vanish as $n\to+\infty$, we find that $p^{*}$ solves the system in Equation (19).
Recall that $p_{n}\in[0,1]^{{\mathcal{X}}}$, $p_{n}(x)=1$ for $x\in A$ and $p_{n}(x)=0$ for $x\in\mathcal{A}_{\mathcal{T}}$ for all $n\in\mathbb{N}$. Hence also $p^{*}\in[0,1]^{{\mathcal{X}}}$, $p^{*}(x)=1$ for $x\in A$ and $p^{*}(x)=0$ for $x\in\mathcal{A}_{\mathcal{T}}$. Due to Proposition 3.9, we can then conclude that $p^{*}$ is the unique minimal solution to the system in Equation (19) and so $p^{*}=\underline{p}^{\mathcal{T}}$. ∎
@@ 4.2 Computing Upper Hitting Probabilities
We show an algorithmic result for computing the upper hitting probability $\overline{p}^{\mathcal{T}}$ similar to the one presented for $\underline{p}^{\mathcal{T}}$. Similarly to lower hitting probabilities, we need to initialise the algorithm with a matrix $T_{1}$ such that $\mathcal{C}_{T_{1}}=\mathscr{W}_{\mathcal{T}}$. We can find such a matrix by taking any point in the interior of $\mathcal{T}$.
@@
Let $T_{1}\in\mathcal{T}$ be any transition matrix such that $\mathcal{C}_{T_{1}}={\mathscr{W}_{\mathcal{T}}}$. For all $n\in\mathbb{N}$, let $p_{n}$ be the minimal nonnegative solution in $0,1]^{{\mathcal{X}}}$ to the linear system
$p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n}p_{n},$
@@
and such that $T_{n+1}(x,\cdot)=T_{n}(x,\cdot)$ if $T_{n}p_{n}](x)=[\overline{T}p_{n}](x)$.Then, the sequence $(p_{n})_{n\in\mathbb{N}}$ is nondecreasing, and its limit $p^{*}\coloneqq\lim_{n\to+\infty}p_{n}$ is the unique minimal solution of the system
$\overline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot\overline{T}\overline{p}^{\mathcal{T}}$ (23)
@@
$\mathcal{C}_{T_{n}}\supseteq\mathcal{C}_{T_{n+1}},$
which implies that $\mathcal{C}_{T_{n}}=\mathscr{W}_{\mathcal{T}}$ for all $n\in\mathbb{N}$.
@@
Let $x\notin\mathcal{C}_{T_{n}}$. We will show that $x\notin\mathcal{C}_{T_{n+1}}$. If $x\in A$, then trivially $x\notin\mathcal{C}_{T_{n+1}}$, so assume $x\in A^{c}$. Then, since $x\notin\mathcal{C}_{T_{n}}$, we have $x\to^{T_{n}}y$ for some $y\in A$. By Lemma 2.14, there exists $m\in\mathbb{N}$ and a sequence of states
@@
On the other hand, if $T_{n+1}p_{n}](x_{k})>[T_{n}p_{n}](x_{k})$ for some $k\in\{0,\ldots,m\}$ let $k^{*}$ be the index of the first such state in the sequence $x_{0},\ldots,x_{m}$. If $x_{k^{*}}\in A$, then, similarly to the previous case, we have $x\to^{T_{n+1}}x_{k^{*}}\in A$ and we are done. Hence, we assume that $x_{k^{*}}\in A^{c}$. Next, define
@@
Let us first assume ex absurdo that $\mathcal{S}=\emptyset$. Since $p_{n}(z)=1\geq p_{n}(x_{k^{*}})$ for $z\in A$, this case occurs only when $p_{n}(x_{k^{*}})=1$. In this case, we find
$1\geq[\overline{T}p_{n}](x_{k^{*}})>[T_{n}p_{n}](x_{k^{*}})=p_{n}(x_{k^{*}})=1,$
where we applied our assumption that
$T_{n+1}p_{n}](x_{k^{*}})>[T_{n}p_{n}](x_{k^{*}}),$
the fact that $p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}}\cdot T_{n}p_{n},$ and our assumption that $x_{k^{*}}\in A^{c}$. Since this yields a contradiction, we have that $\mathcal{S}\neq\emptyset$.

### v2

@@ 3.3 Upper Hitting Probabilities
Let $q\in[0,1]^{\mathcal{X}}$ be a solution of (18). By minimality, we know that $q\geq\overline{p}^{\mathcal{T}}$, therefore we need to show $q\leq\overline{p}^{\mathcal{T}}$. Since $q$ and $\overline{p}^{\mathcal{T}}$ already agree on $A$ and $\mathscr{W}_{\mathcal{T}}$ it suffices to show that $q|_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}\leq\overline{p}^{\mathcal{T}}|_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}$. By Proposition 2.16, there exists $S\in\mathcal{T}$ such that $\overline{T}q=Sq$. It follows that $q=\mathbb{1}_{A}+\mathbb{1}_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}\cdot Sq.$ From (18), we obtain
$\overline{p}^{\mathcal{T}}|_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}=(\overline{T}\overline{p}^{\mathcal{T}})|_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}\geq(S\overline{p}^{\mathcal{T}})|_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}.$
Now we apply Lemma 2.6 and we get
$\overline{p}^{\mathcal{T}}|_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}\geq(S\overline{p}^{\mathcal{T}})|_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}=S|_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}\overline{p}^{\mathcal{T}}|_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}+(S\mathbb{1}_{A})|_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}.$
Recursively expanding this expression for $\overline{p}^{\mathcal{T}}|_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}$ yields, after $m\in\mathbb{N}$ expansions
$\overline{p}^{\mathcal{T}}|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}\geq(S|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}})^{m}\ \overline{p}^{\mathcal{T}}|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}+\left[\sum_{k=0}^{m-1}(S|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}})^{k}\right](S\mathbb{1}_{A})|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}.$
Since $\mathcal{C}_{S}={\mathscr{W}_{\mathcal{T}}}$, we let $m\to+\infty$ and we apply Lemma 2.12:
$\overline{p}^{\mathcal{T}}|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}\geq(I-S|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}})^{-1}(S\mathbb{1}_{A})|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}}=q|_{A^{c}\setminus{\mathscr{W}_{\mathcal{T}}}},$
where the last equality follows from Proposition 2.13. Thus $q=\overline{p}^{\mathcal{T}}$, which is what we wanted to prove. ∎
@@ 3.4 Further Discussions on the Definition of Reachability
Proposition 3.11 shows that this definition is equivalent to $\operatorname{(LR0)}$: indeed, $x\in\mathcal{A}_{\mathcal{T}}$, i.e. $x\not\rightharpoondown A$, if and only if $\underline{p}^{\mathcal{T}}(x)=0$. Taking the negation, we obtain that $x$ lower reaches $A$ under (LR0) if and only if $\underline{p}^{\mathcal{T}}(x)>0$, thus definitions $\operatorname{(LR0)}$ and $\operatorname{(LRHP)}$ are equivalent.
In the same way, the definition of upper reachability can be rewritten in terms of positivity of the upper hitting probability. These definitions of reachability for IMCs mirror the behaviour in the precise setting. In fact, also for a (precise) homogeneous Markov chain we have that $x$ reaches $A$ if and only if its hitting probability is positive, i.e. $p^{T}(x)>0$.
@@ 4.1 Computing Lower Hitting Probabilities
We show an algorithmic result for computing $\underline{p}^{\mathcal{T}}$. This result, which emulates Krak [20, Proposition 7], provides a practical algorithm for calculating $\underline{p}^{\mathcal{T}}$ by iterating over extreme points of $\mathcal{T}$. The only condition is to start with $T_{1}\in\mathcal{T}$ such that the trivial hitting probabilities are already correct (i.e. $p_{1}(x)=0$ for $x\in\mathcal{A}_{\mathcal{T}}$ and $p_{1}(x)=1$ for $x\in A$). Definition (LR1) provides a practical algorithm to identify the set $\mathcal{A}_{\mathcal{T}}$ and the transition matrix $T_{1}\in\mathcal{T}$. In particular, if we set $\mathcal{D}=A$ then $\mathcal{D}_{k^{*}}^{c}=\mathcal{A}_{\mathcal{T}}$. For all $z\in\mathcal{A}_{\mathcal{T}}=\mathcal{D}_{k^{*}}^{c}$ we have $\underline{T}\mathbb{1}_{\mathcal{D}_{k^{*}}}](z)=0$, therefore there exists a matrix $S_{z}\in\mathcal{T}$ such that $S_{z}\mathbb{1}_{\mathcal{D}_{k^{*}}}](z)=0$. We construct the initial matrix $T_{1}\in\mathcal{T}$ by setting $T_{1}(z,\cdot)=S_{z}(z,\cdot)$, for all $z\in\mathcal{A}_{\mathcal{T}}$, and an arbitrary admissible row for any other state. It follows that $\mathcal{C}_{T_{1}}=\mathcal{A}_{\mathcal{T}}$.
@@
Let $T_{1}\in\mathcal{T}$ be any transition matrix such that $\mathcal{C}_{T_{1}}=\mathcal{A}_{\mathcal{T}}$. For all $n\in\mathbb{N}$, let $p_{n}$ be the vector of hitting probabilities for $T_{n}$, i.e. the unique solution of the linear system
$p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}\setminus\mathcal{C}_{T_{n}}}\cdot T_{n}p_{n},$
@@
Then, the sequence $(p_{n})_{n\in\mathbb{N}}$ is nonincreasing and its limit $p^{*}\coloneqq\lim_{n\to+\infty}p_{n}$ is the vector of lower hitting probabilities, i.e. the unique solution of the system
$\underline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}\setminus\mathcal{A}_{\mathcal{T}}}\cdot\underline{T}\underline{p}^{\mathcal{T}}.$ (19)
@@
The proof uses induction on (20). For the base case $n=1$, we have $\mathcal{C}_{T_{1}}=\mathcal{A}_{\mathcal{T}}$ by the choice of $T_{1}\in\mathcal{T}$. Assume now that $\mathcal{C}_{T_{n}}=\mathcal{A}_{\mathcal{T}}$ for some $n\in\mathbb{N}$. We show that also $\mathcal{C}_{T_{n+1}}=\mathcal{A}_{\mathcal{T}}$. First, we observe that $\mathcal{C}_{T_{n+1}}\subseteq\mathcal{A}_{\mathcal{T}}=\bigcup_{T\in\mathcal{T}}\mathcal{C}_{T}$, by Lemma 3.10. What remains to prove is $\mathcal{A}_{\mathcal{T}}\subseteq\mathcal{C}_{T_{n+1}}$. Let $x\in\mathcal{A}_{\mathcal{T}}$. Since $p_{n}$ is the vector of hitting probabilities for $T_{n}$, by Proposition 2.5 we have
$T_{n}p_{n}](x)=p_{n}(x)=0,$
where the last equality follows from the induction hypothesis $\mathcal{C}_{T_{n}}=\mathcal{A}_{\mathcal{T}}$. By construction, we have
$T_{n+1}p_{n}=\underline{T}p_{n}\leq T_{n}p_{n}.$
Therefore,
$0=[T_{n}p_{n}](x)\geq[T_{n+1}p_{n}](x)=\sum_{y\in\mathcal{X}}T_{n+1}(x,y)p_{n}(y)\geq 0.$
Since the right-hand side is nonnegative, we must have $T_{n+1}(x,y)=0$ for every $y\in\mathcal{X}$ with $p_{n}(y)>0$. By Proposition 2.5 this implies $T_{n+1}(x,y)=0$ for all $y\notin\mathcal{C}_{T_{n}}=\mathcal{A}_{\mathcal{T}}$. From the arbitrariness of $x\in\mathcal{A}_{\mathcal{T}}$ we have $T_{n+1}(x,y)=0$ for all $x\in\mathcal{A}_{\mathcal{T}}$ and $y\notin\mathcal{A}_{\mathcal{T}}$. Since $A\subseteq\mathcal{A}_{\mathcal{T}}^{c}$, we have $x\not\to^{T_{n+1}}A$ for all $x\in\mathcal{A}_{\mathcal{T}}$. Thus $x\in\mathcal{C}_{T_{n+1}}$ and $\mathcal{A}_{\mathcal{T}}\subseteq\mathcal{C}_{T_{n+1}}$, concluding the proof. ∎
@@
The sequence $(p_{n})_{n\in\mathbb{N}}$ is well-defined since
$p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}\setminus\mathcal{C}_{T_{n}}}\cdot T_{n}p_{n}$
has a unique solution (Proposition 2.13) and the existence of the extreme point $T_{n+1}$ is guaranteed by Proposition 2.16.
By Proposition 2.13, we have that, for all $n\in\mathbb{N}$, $p_{n}(x)=1$ for all $x\in A$ and $p_{n}(x)=0$ for all $x\in\mathcal{C}_{T_{n}}=\mathcal{A}_{\mathcal{T}}$, therefore $(p_{n})_{n\in\mathbb{N}}$ is trivially convergent on $A\cup\mathcal{A}_{\mathcal{T}}$.
Next, we show that $(p_{n})_{n\in\mathbb{N}}$ is nonincreasing and convergent on $A^{c}\setminus\mathcal{A}_{\mathcal{T}}$. Recall that $T_{n+1}p_{n}=\underline{T}p_{n}\leq T_{n}p_{n}$ for all $n\in\mathbb{N}$. This yields
@@
where we used Lemma 2.6 and the fact that $\mathcal{A}_{\mathcal{T}}=\mathcal{C}_{T_{n}}$ for all $n\in\mathbb{N}$. Repeatedly expanding this inequality $m\in\mathbb{N}$ times yields
@@
Letting $m\to+\infty$ and applying Lemma 2.12 yields
@@
where the last step follows from Proposition 2.13. Thus the sequence $(p_{n})_{n\in\mathbb{N}}$ is nonincreasing. Since $(p_{n})_{n\in\mathbb{N}}$ is nonincreasing and $p_{n}\in[0,1]^{{\mathcal{X}}}$ for all $n\in\mathbb{N}$, it follows that the limit $p^{*}\coloneqq\lim_{n\to+\infty}p_{n}$ exists.
We move on to showing that $p^{*}$ satisfies the system in Equation (19). Fix any $n\in\mathbb{N}$. Then
@@
[…]
for any vector norm $||\cdot||$. Continuing (21)
@@
[…]
where we used Lemma 2.17(3) for the second and third inequality. Since all summands on the final right-hand side of (22) vanish as $n\to+\infty$, we find that $p^{*}$ is a solution of the system in Equation (19). Since Proposition 3.11 guarantees uniqueness of the solution, it follows that $p^{*}=\underline{p}^{\mathcal{T}}$. ∎
@@ 4.2 Computing Upper Hitting Probabilities
We present an algorithm to compute the upper hitting probability $\overline{p}^{\mathcal{T}}$, analogous to the one for $\underline{p}^{\mathcal{T}}$. As with lower hitting probabilities, the algorithm must be initialised with a matrix $T_{1}$ satisfying $\mathcal{C}_{T_{1}}=\mathscr{W}_{\mathcal{T}}$. Any interior point $T_{1}\in\mathcal{T}$ satisfies this condition because $T_{1}(x,y)>0$ if and only if $\overline{T}\mathbb{1}_{\{y\}}](x)>0$.
@@
Let $T_{1}\in\mathcal{T}$ be any transition matrix such that $\mathcal{C}_{T_{1}}={\mathscr{W}_{\mathcal{T}}}$. For all $n\in\mathbb{N}$, let $p_{n}$ be the vector of hitting probabilities for $T_{n}$, i.e. the unique solution of the linear system
$p_{n}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}\setminus\mathcal{C}_{T_{n}}}\cdot T_{n}p_{n},$
@@
and such that $T_{n+1}(x,\cdot)=T_{n}(x,\cdot)$ if $T_{n}p_{n}](x)=[\overline{T}p_{n}](x)$. Then, the sequence $(p_{n})_{n\in\mathbb{N}}$ is nondecreasing, and its limit $p^{*}\coloneqq\lim_{n\to+\infty}p_{n}$ is the vector of upper hitting probabilities, i.e the minimal nonnegative solution of the system
$\overline{p}^{\mathcal{T}}=\mathbb{1}_{A}+\mathbb{1}_{A^{c}\setminus\mathscr{W}_{\mathcal{T}}}\cdot\overline{T}\overline{p}^{\mathcal{T}}.$ (23)
@@
$\mathcal{C}_{T_{n}}=\mathscr{W}_{\mathcal{T}}$
for all $n\in\mathbb{N}$. Equivalently, $\mathcal{C}_{T_{n}}\supseteq\mathcal{C}_{T_{n+1}}$.
@@
Let $x\notin\mathcal{C}_{T_{n}}$. We show that $x\notin\mathcal{C}_{T_{n+1}}$. If $x\in A$, then trivially $x\notin\mathcal{C}_{T_{n+1}}$, so assume $x\in A^{c}$. Then, since $x\notin\mathcal{C}_{T_{n}}$, we have $x\to^{T_{n}}y$ for some $y\in A$. By Lemma 2.14, there exists $m\in\mathbb{N}$ and a sequence of states
@@
If $T_{n+1}p_{n}](x_{k})>[T_{n}p_{n}](x_{k})$ for some $k\in\{0,\ldots,m\}$, let $k^{*}$ be the index of the first such state in the sequence $x_{0},\ldots,x_{m}$. If $x_{k^{*}}\in A$, then, similarly to the previous case, we have $x\to^{T_{n+1}}x_{k^{*}}\in A$ and we are done. Hence, we assume that $x_{k^{*}}\in A^{c}$. Define
@@
Let us assume ex absurdo that $\mathcal{S}=\emptyset$. Since $p_{n}(z)=1\geq p_{n}(x_{k^{*}})$ for $z\in A$, $\mathcal{S}=\emptyset$ occurs only if $p_{n}(x_{k^{*}})=1$. In this case, we find
$1\geq[T_{n+1}p_{n}](x_{k^{*}})>[T_{n}p_{n}](x_{k^{*}})=p_{n}(x_{k^{*}})=1,$
which is a contradiction, so $\mathcal{S}\neq\emptyset$.